Analysis of a sample of an organic compound showed it to contain 39.9% carbon 6.9% hydrogen and 53.3% oxygen


1
Expert's answer
2021-01-11T03:46:41-0500

There is no question itself, but this data is likely given to determine the empirical formula of the compound.

100 grams of this compound contain 39.9 g (C), 6.9 g (H) and 53.3 g (O).

Amounts of each element in moles are:

"n(C)=\frac{39.9g}{12.01g/mol}=3.32mol"


"n(H)=\frac{6.9g}{1.01g/mol}=6.8mol"


"n(O)=\frac{53.3g}{16.0g/mol}=3.33mol"


The C : H : O molar ratio appears to be 1 : 2 : 1, hence the empirical formula is CH2O.

The molecular formula cannot be determined without knowing the molar mass of the compound.